# Auslander-Reiten translations in the monomorphism categories of exact categories

### v1

@@ 2.2. Almost split sequences
    For each indecomposable non-injective object $A$, there is an almost split sequence $A\to B\to C$.
Let ${\mathcal{C}}$ be a Hom-finite $k$-linear Krull-Schmidt category. Denote by ${\mathcal{C}}\textendash\operatorname{Mod}$ the category of additive contravariant functors from ${\mathcal{C}}$ to abelian groups and ${\mathcal{C}}\textendash\operatorname{mod}$ the subcategory consisting of finitely presented functors. Then ${\mathcal{C}}$ is called a dualizing $k$-variety [AR74] if the $k$-dual functors $D:{\mathcal{C}}\textendash\operatorname{Mod}\to{\mathcal{C}}^{\rm op}\textendash\operatorname{Mod}$ and $D:{\mathcal{C}}^{\rm op}\textendash\operatorname{Mod}\to{\mathcal{C}}\textendash\operatorname{Mod}$ given by $D(F)(C)=D(FC)$ for every object $C$ of ${\mathcal{C}}$ and $F\in{\mathcal{C}}\textendash\operatorname{Mod}$ or ${\mathcal{C}}^{\rm op}\textendash\operatorname{Mod}$ induce dualities
$D:{\mathcal{C}}\textendash\operatorname{mod}\to{\mathcal{C}}^{\rm op}\textendash\operatorname{mod}\ \mbox{ and }\ D:{\mathcal{C}}^{\rm op}\textendash\operatorname{mod}\to{\mathcal{C}}\textendash\operatorname{mod}.$
In this case, it turns out that ${\mathcal{C}}\textendash\operatorname{mod}$ is an abelian subcategory of ${\mathcal{C}}\textendash\operatorname{Mod}$ that admits enough projective and enough injective objects [AR74, Theorem 2.4]. We also note that if ${\mathcal{C}}$ is a dualizing $k$-variety then so is the functor category ${\mathcal{C}}\textendash\operatorname{mod}$ ([AR74, Proposition 2.6]). In particular, ${\mathcal{C}}\textendash\operatorname{mod}$ is also Krull-Schmidt. Furthermore, any functorially finite subcategory of a dualizing $k$-variety is a dualizing $k$-variety by [AS, Theorem 2.3].
@@
[INP, Corollary 4.5] Let ${\mathcal{C}}$ be a $k$-linear $\operatorname{Ext}$-finite Krull-Schmidt exact category with enough projectives and injectives. Then the following conditions are equivalent.

### v2

@@ 3.3. Almost split sequences in morphism categories
    The following exact sequences are almost split in ${{\mathcal{F}}}({\mathcal{C}})$.
    ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\left(\begin{smallmatrix}A\\ 0\end{smallmatrix}\right)}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}f\\ 0\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}B\\ C\end{smallmatrix}\right)}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}g\\ 1\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}C\\ C\end{smallmatrix}\right)}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
    ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\left(\begin{smallmatrix}A\\ A\end{smallmatrix}\right)}_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}[1,0]^{T}\\ f\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}A\oplus P(C)\\ B\end{smallmatrix}\right)}_{[f,p^{\prime}]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}[0,1]\\ g\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}P(C)\\ C\end{smallmatrix}\right)}_{p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
Proof.
As the proof is essentially the same as in the module category, we refer to [RS2, Proposition 7.4]. ∎
Next, we show that non-sporadic almost split sequences in ${\mathcal{H}}({\mathcal{C}})$ (resp. ${{\mathcal{S}}}({\mathcal{C}})$, ${{\mathcal{F}}}({\mathcal{C}})$) are cw-exact.
Lemma 3.9.
Let $\mathcal{C}$} be an exact category and
${\delta:0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)}_{f}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\phi_{1}\\ \phi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Z_{1}\\ Z_{2}\end{smallmatrix}\right)}_{h}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\left(\begin{smallmatrix}\psi_{1}\\ \psi_{2}\end{smallmatrix}\right)}$${{\left(\begin{smallmatrix}Y_{1}\\ Y_{2}\end{smallmatrix}\right)}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
be an almost split sequence in ${{\mathcal{H}}}({\mathcal{C}})$ (resp. ${{\mathcal{S}}}({\mathcal{C}})$, ${{\mathcal{F}}}({\mathcal{C}})$). If $\delta$ is not isomorphic to one of the almost split sequences in Lemma 3.8 (1) (resp. (2), (3)), then the sequences $0\rightarrow X_{i}\stackrel{{\scriptstyle\phi_{i}}}{{\rightarrow}}Z_{i}\stackrel{{\scriptstyle\psi_{i}}}{{\rightarrow}}Y_{i}\rightarrow 0$, $i=1,2$, split.
Proof.
We just prove for the case $\delta\in{\mathcal{H}}({\mathcal{C}})$. Since $\delta$ is not isomorphic to one of the sequences in Lemma 3.8 $(1)$, $g$ is not a split monomorphism. It is easy to check that $\left(\begin{smallmatrix}1\\ g\end{smallmatrix}\right):\left(\begin{smallmatrix}Y_{1}\\ Y_{1}\end{smallmatrix}\right)_{1}\to{\left(\begin{smallmatrix}Y_{1}\\ Y_{2}\end{smallmatrix}\right)}_{g}$ is not a split epimorphism. Hence $\left(\begin{smallmatrix}1\\ g\end{smallmatrix}\right)$ factors through ${\left(\begin{smallmatrix}Z_{1}\\ Z_{2}\end{smallmatrix}\right)}_{h}$ i.e. the top row splits. Similarly, $f$ is not a split epimorphism. Thus $\left(\begin{smallmatrix}f\\ 1\end{smallmatrix}\right):\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)_{f}\to{\left(\begin{smallmatrix}X_{2}\\ X_{2}\end{smallmatrix}\right)}_{1}$ is not a split monomorphism and hence factors through ${\left(\begin{smallmatrix}Z_{1}\\ Z_{2}\end{smallmatrix}\right)}_{h}$. So the bottom row splits.
A similar argument as [RS2, Corollary 7.6] shows that if $\delta$ is an almost split sequences in ${\mathcal{S}}({\mathcal{C}})$ or ${\mathcal{F}}({\mathcal{C}})$, which is not in sporadic cases (2) and (3), then both rows of $\delta$ split. ∎
Now we discuss the key technique, which allows us to transfer cw-exact almost split sequences in ${\mathcal{H}}({\mathcal{C}})$ to almost split sequences in ${\mathcal{C}}\textendash\operatorname{mod}$.
Construction 3.10.
[HE, Construction 4.1] Let $(X_{1}\stackrel{{\scriptstyle f}}{{\rightarrow}}X_{2})$ be an object of ${{\mathcal{H}}}({\mathcal{C}})$. Define
$(X_{1}\stackrel{{\scriptstyle f}}{{\rightarrow}}X_{2})\stackrel{{\scriptstyle\Theta}}{{\mapsto}}\ \operatorname{Coker}\operatorname{Hom}_{{\mathcal{C}}}(-,f).$
If $\phi=\left(\begin{smallmatrix}\phi_{1}\\ \phi_{2}\end{smallmatrix}\right):X=\left(\begin{smallmatrix}X_{1}\\ X_{2}\end{smallmatrix}\right)_{f}\rightarrow\left(\begin{smallmatrix}Z_{1}\\ Z_{2}\end{smallmatrix}\right)_{h}=Z$ is a morphism in ${\mathcal{H}}({\mathcal{C}})$, $\Theta(\phi)$ is defined to be the unique morphism that makes the bottom left square of the following Figure 1 commutes.
Let $\mathrm{Y}=(Y_{1}\stackrel{{\scriptstyle g}}{{\rightarrow}}Y_{2})$ be an indecomposable object in ${\mathcal{H}}({\mathcal{C}})$ which is neither a split monomorphism nor a split epimorphism. Take
[…]
to be the almost split sequence in ${\mathcal{H}}({\mathcal{C}})^{{\rm cw}}$ ending at $\mathrm{Y}$. For simplicity, set $\mathrm{Z}=(Z_{1}\stackrel{{\scriptstyle h}}{{\rightarrow}}Z_{2})$, $\mathrm{X}=(X_{1}\stackrel{{\scriptstyle f}}{{\rightarrow}}X_{2})$, $\phi=(\phi_{1},\phi_{2})^{T}$ and $\psi=(\psi_{1},\psi_{2})^{T}$.
Assume $\mathcal{C}$} has pseudokernel. By applying $\Theta$, one gets the commutative diagram with exact rows.
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(-,X_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-,f)}$$\scriptstyle{(-,\phi_{1})}$${(-,Z_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-,h)}$$\scriptstyle{(-,\psi_{1})}$${(-,Y_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-,g)}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(-,X_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-,\phi_{2})}$${(-,Z_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(-,\psi_{2})}$${(-,Y_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${\Theta(\mathrm{X})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Theta(\phi)}$${\Theta(\mathrm{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Theta(\psi)}$${\Theta(\mathrm{Y})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0}$${0}$${0}$
Figure 1.
in ${\mathcal{C}}\textendash\operatorname{mod}$ whose bottom row is indeed the image $\Theta(\delta)$ of $\delta$ under the functor $\Theta$.
@@
[HE, Lemma 4.8, Theorem 4.9] Let $\mathcal{C}$} be an additive category having peudokernels. In the above diagram,
  1. (1)
    the map $\Theta(\phi)$ is a monomorphism. Thus, $\Theta(\delta)$ is a short exact sequence in ${\mathcal{C}}\textendash\operatorname{mod}.$
  2. (2)
    $\Theta(\delta)$ is an almost split sequence in ${\mathcal{C}}\textendash\operatorname{mod}$.
3.4. Auslander-Reiten translation in functor categories
Let $\mathcal{V}$} be a Hom-finite $k$-linear Krull-Schmidt category. Denote by ${\mathcal{V}}\textendash\operatorname{Mod}$ the category of additive contravariant functors from $\mathcal{V}$} to abelian groups and ${\mathcal{V}}\textendash\operatorname{mod}$ the subcategory consisting of finitely presented functors. Then $\mathcal{V}$} is called a dualizing $k$-variety [AR74] if the $k$-dual functors $D:{\mathcal{V}}\textendash\operatorname{Mod}\to{\mathcal{V}}^{\rm op}\textendash\operatorname{Mod}$ and $D:{\mathcal{V}}^{\rm op}\textendash\operatorname{Mod}\to{\mathcal{V}}\textendash\operatorname{Mod}$ given by $D(F)(C)=D(FC)$ for every object $C$ of $\mathcal{V}$} and $F\in{\mathcal{V}}\textendash\operatorname{Mod}$ or ${\mathcal{V}}^{\rm op}\textendash\operatorname{Mod}$ induce dualities
[…]
Let $\mathcal{V}$} be a dualizing $k$-variety. It turns out that ${\mathcal{V}}\textendash\operatorname{mod}$ is an abelian subcategory of ${\mathcal{V}}\textendash\operatorname{Mod}$ that admits enough projective and enough injective objects [AR74, Theorem 2.4]. The following result shows a connection among almost split sequences, Auslander-Reiten duality, and dualizing vareities.
Theorem 3.12.
[INP, Corollary 4.5] Let $\mathcal{C}$} be a $k$-linear $\operatorname{Ext}$-finite Krull-Schmidt exact category with enough $\operatorname{Ext}$-projectives and $\operatorname{Ext}$-injectives. Then the following conditions are equivalent.